OpenStudy (anonymous):

What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i want to know that question too

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

If your angle measure is in degrees, then the area of a sector is given by \[\huge \frac{\theta^o}{360^o}\pi r^2\]